• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E. Cisbani

    Log entry time 08:03:14 on June26,2005

    Entry number 147254

    keyword=Owl Shift summary

    Owl Shift summary

    No beam for the first two hours of the shift

    04:25 Beam off, problem with couple of vacuum valves in HallA; Access required. Couple of hours to figure out the problem and find the expert. No beam available for the rest of the shift.

    Runs Summary

    01:55-02:37 Run 4914 (H2O speed = 601.3)
    02:37-03:17 Run 4915 (H2O speed = 602.5)
    03:18-03:55 Run 4916 (H2O speed = 603.8)
    03:55-04:25 Run 4917 (H2O speed = 602.9) Beam off